OpenStudy (anonymous):

4 questions please help: 1. x^2 + 13x + 40/ x-7 divided by x+8/x^2 - 49 2. (12x^2 + 9x - 7) divided by 3x 3. (3d^2 + 2d - 29 divided (d+3) 4. What is the length of the rectangle when the width and area is given? A = (4b^3 + 5b - 3) in^2. W = (2b - 1) in.

OpenStudy (campbell_st):

well the rule for dividing with fractions is take the reciprocal of the denominator and multiply...(flip and multiply) so you have \[\frac{x^2 + 13x + 40}{x - 7} \times \frac{x^2 - 49}{x + 8}\] if you factorise the quadratics you'll see some common factors that can be eliminated.